OpenStudy (anonymous):

Which of the following is true about the function below? 1/√ x-6 A. Its domain is [6,∞) and its range is . (-∞,0) U (0,∞) B. Its domain is (6,∞) and its range is . (0,∞) C. Its domain is (6,∞) and its range is . (-∞,∞) D. Its domain is (-∞,0] and its range is . (0,∞)

OpenStudy (anonymous):

The domain will be all numbers that prevent the denominator from becoming zero, or otherwise undefined. Where is \(\sqrt{x-6}\) defined? In other words, finding the domain of this will give the domain of \(\dfrac{1}{\sqrt{x-6}}\). As for range, consider what kind of values can't be attained. Obviously, you can get negative numbers because the square root function never yields a negative number, so you can eliminate that part of the potential range, \((-\infty,0)\). What about 0? Can the fraction ever attain 0? What about the positive numbers?

OpenStudy (anonymous):

so everyone that has a zero is wrong?

OpenStudy (anonymous):

I can't draw an answer from what you gave me.

OpenStudy (anonymous):

For the fraction to be defined, you must have \(x>6\). If \(x=6\), you get \(\dfrac{1}{\sqrt0}\), which is undefined. If \(x<6\), you get \(\dfrac{1}{\sqrt{\text{negative number}}}\), also undefined. So, the domain is all numbers strictly greater than 6, or \((6,\infty)\).

OpenStudy (anonymous):

For the range, I suggested you think about what kind of values the fraction can't possibly attain. Can \(\dfrac{1}{\sqrt{\cdots}}\) ever be negative? (No, the square root only yields positive values.) Can \(\dfrac{1}{\sqrt{\cdots}}\) ever be zero? (No, this is not possible for any fraction so long as the numerator is non-zero. Think about \(\dfrac{1}{x}\), which is undefined at 0, as an example.) This means the range must be \((0,\infty)\), since nothing is barring the fraction from taking on any positive number.